# Large Language Model Compression with Global Rank and Sparsity Optimization

### v1

@@ 5.2.2 Learnable Probabilistic Pruning
$\tilde{\mathbf{W}}\;=\;\mathbf{U}\,\mathrm{diag}\!\bigl{(}\boldsymbol{\sigma}\odot\mathbf{m}_{\sigma}\bigr{)}\,\mathbf{V}^{\top}\;+\;\mathbf{S}\,\odot\,\mathbf{m}_{S},$ (12)
@@
$\min_{\mathbf{s}}\ \mathbb{E}_{\mathbf{m}\sim p(\mathbf{m}\mid\mathbf{s})}\Bigl{[}\,\mathcal{L}(\tilde{\mathbf{W}})\,\Bigr{]},$ (13)
where $\mathbf{s}=\{s_{\sigma_{i}},s_{S_{ij}}\}$ and $p(\mathbf{m}\mid\mathbf{s})$ is the product of Bernoulli distributions. We employ a REINFORCE-style Williams [1992] policy gradient:
@@
with a small $\epsilon>0$ to avoid division by zero. To reduce variance, we maintain a moving average baseline $\delta$ Zhao et al. [2011]:
$\delta\;\leftarrow\;\beta\,\delta\;+\;(1-\beta)\,\mathcal{L}(\tilde{\mathbf{W}}),$ (15)
@@
Thresholding masks and final factorization.
Once the probabilities converge, we determine the binary masks to satisfy the desired compression ratio. Let the total number of parameters be $P_{\text{total}}$, and the target compression ratio be $r$. We aim to retain $P_{\text{keep}}=(1-r)\cdot P_{\text{total}}$ parameters.
Each singular value $\sigma_{i}$ contributes $(m+n)$ parameters, and each nonzero entry of $\mathbf{S}$ contributes 1 parameter. To determine which parameters to retain, we compute a weighted score for all potential parameters as the product of their retention probability $s_{k}$ and their contribution. These scores are sorted in descending order, and the top $P_{\text{keep}}$ parameters are retained, generating binary masks $m_{k}$:
$m_{k}\;=\;\begin{cases}1,&\text{if }k\text{ is among the top }P_{\text{keep}},\\ 0,&\text{otherwise}.\end{cases}$ (17)
We reconstruct the compressed weight matrix using Eq. (12). Specifically, the low-rank component $\mathbf{L}$ is reconstructed by selecting the singular vectors and values corresponding to the retained singular values. To further reduce the storage and computational cost, we factorize the low-rank component into smaller matrices.
The compressed $\mathbf{U}^{\prime}$ and $\mathbf{V}^{\prime}$ matrices are computed as:
@@
where $r^{\prime}$ is the number of retained singular values (i.e., $m_{\sigma_{i}}=1$), and $\mathbf{u}_{i}$, $\mathbf{v}_{i}$ are the corresponding singular vectors.
The final compressed weight matrix is then:
@@
This factorization reduces both the storage requirements and the computational cost during inference, as matrix multiplications involving $\mathbf{W}$ can be replaced with smaller matrices $\mathbf{U}^{\prime}$ and $\mathbf{V}^{\prime}$.
@@ 5.3 Discussion
Our method introduces a two-stage approach for compressing LLMs, which effectively addresses key challenges in model compression. In the first stage, we apply RPCA to decompose the weight matrix into low-rank and sparse components. This decomposition reduces the optimization space significantly, allowing us to focus on the most important components. By separating global structures from local anomalies, RPCA addresses the first major challenge in model compression. In the second stage, we further refine the model by optimizing the retention probabilities of both the low-rank and sparse components. This fine selection process helps detect and handle redundancy across different layers, allowing us to efficiently manage layer-wise redundancy. Through this two-stage process, our method adapts dynamically to the model’s structure and optimally prunes unnecessary parameters.
@@ 6 Experiments
Model w/o Pruning SparseGPT WANDA BESA Our Method
LLaMA-7B 5.68 66.31 7.22 63.12 7.26 61.81 6.86 63.13 6.61 64.29
LLaMA-13B 5.09 68.91 6.21 65.98 6.15 66.49 5.92 67.43 5.76 68.32
LLaMA-30B 4.10 72.73 5.33 70.53 5.25 70.92 5.00 71.61 4.77 72.08
LLaMA-2 7B 5.21 66.96 6.99 63.71 6.92 63.81 6.60 64.92 6.25 65.33
LLaMA-2 13B 4.88 69.95 6.02 67.22 5.97 67.94 5.75 68.45 5.49 69.14
Figure 3: Perplexity changes with sequential decomposition and pruning of each layer in LLaMA2-7B.
In this section, we first introduce the experimental setup, including datasets, evaluation metrics, implementation details, and baseline methods. Subsequently, we present the main experimental results in §6.1. Due to space constraints, comparisons under different compression rates can be found in Appendix D, while the comparison with LoSparse is provided in Appendix B, and the comparison with LPAF is available in Appendix E. Finally, in §6.2, we conduct ablation studies and further discussions.
@@
We evaluate our proposed method on widely adopted large language models, including LLaMA 7B/13B/30B Touvron et al. [2023a] and LLaMA-2 7B/13B Touvron et al. [2023b], as well as BERT-base Devlin et al. [2018] and DeBERTaV3-base He et al. [2023]. To assess the performance of the compressed models, we conduct experiments on zero-shot tasks and language modeling. In addition to measuring perplexity, we perform an extensive evaluation of the zero-shot capabilities of pruned models across six standard commonsense benchmark datasets. These benchmarks include PIQA Bisk et al. [2020], BoolQ Clark et al. [2019], HellaSwag Zellers et al. [2019], WinoGrande Sakaguchi et al. [2021], and the ARC Easy and ARC Challenge tasks Clark et al. [2018]. Following previous work on LLM compression Xiao et al. [2023], Frantar and Alistarh [2023], we evaluate perplexity on the held-out WikiText Merity et al. [2016] validation set.
@@
We utilize the following configurations: PyTorch version 2.3.0, Transformers library version 4.28.0, CUDA version 12.1, GPU: NVIDIA A100. Operating System: Ubuntu. To ensure a fair comparison, we use the same set of calibration data across methods when required. Specifically, we use 128 sequences with context length sampled from the C4 training set Raffel et al. [2020], as used in prior works. For the policy gradient estimation, we set the number of iterations to 3, with a sliding window size of 5 and a learning rate of 0.05. These configurations were selected to balance computational efficiency and optimization stability.

### v3

@@ 2.3.2 Stage 2: Learnable Probabilistic Pruning
$\tilde{\mathbf{W}}\;=\;\mathbf{U}\,\mathrm{diag}\!\bigl(\bm{\sigma}\odot\mathbf{m}_{\sigma}\bigr)\,\mathbf{V}^{\top}\;+\;\mathbf{S}\,\odot\,\mathbf{m}_{S},$ (9)
@@
[…]
where $\mathbf{s}=\{s_{\sigma_{i}},s_{S_{ij}}\}$ and $p(\mathbf{m}\mid\mathbf{s})$ is the product of Bernoulli distributions. We employ a REINFORCE-style (Williams, 1992) policy gradient:
@@
with a small $\epsilon>0$ to avoid division by zero. To reduce variance, we maintain a moving average baseline $\delta$ (Zhao et al., 2011):
[…]
@@
After each gradient step, we project $\mathbf{s}$ back onto $\{\mathbf{s}:\mathbf{1}^{\top}\mathbf{s}\leq K,\;0\leq s_{k}\leq 1\}$.
Thresholding masks and final factorization.
The policy gradient optimization yields a set of probabilities $\{s_{k}\}$ that reflect the learned importance of each parameter for minimizing the task loss. To obtain the final compressed model that strictly adheres to the budget $K$, we perform a deterministic selection. We treat the probabilities $s_{k}$ as importance scores for their corresponding parameters (singular values or sparse entries). All potential parameters are ranked globally according to these scores. We select the top-$K$ parameters to keep, generating the final binary masks $m_{k}$:
@@
We note that comparing $s_{k}$ across different parameter types is valid because the learned probability acts as a unified proxy for the “utility-to-cost” ratio. During optimization, the policy gradient inherently accounts for the parameter’s contribution to loss reduction relative to its existence in the model. Thus, sorting by $s_{k}$ provides a globally consistent ranking for resource allocation.
This final step ensures the parameter budget is met precisely. The compressed weight matrix is reconstructed using these binary masks in Eq. equation 9. To enhance efficiency, the resulting low-rank component is factorized into smaller matrices. The compressed $\mathbf{U}^{\prime}$ and $\mathbf{V}^{\prime}$ are computed as:
@@
where $r^{\prime}$ is the number of retained singular values (i.e., where $m_{\sigma_{i}}=1$). The final compressed weight matrix is then:
@@
This factorization reduces both storage and computational cost during inference.
@@ 2.4 Discussion
We propose CAP, a two-stage compression framework for large language models. Stage 1: RPCA Decomposition—The weight matrix is split into low-rank and sparse parts, preserving global structure while isolating local anomalies and sharply reducing the search space for later optimization. This step is cast as a convex program (nuclear norm + $L_{1}$ norm), guaranteeing a globally optimal separation for the decomposition objective. Stage 2: Bernoulli Mask Optimization—Using a small calibration set, an unbiased policy-gradient method learns the retention probabilities for the low-rank and sparse components, automatically detecting and pruning redundancy across layers.
Remark. While Stage 1 solves a convex problem, Stage 2 addresses the discrete budget allocation problem via policy gradient, which serves as a heuristic optimizer. It does not guarantee a global optimum for the non-convex pruning objective but empirically finds effective allocation strategies by leveraging the high-quality subspaces identified in Stage 1. This two-step design balances theoretical principledness in decomposition with practical flexibility in resource allocation.
Table 1: Performance comparison with unstructured pruning methods at 50% compression. We report average zero-shot accuracy (%) across eight tasks and WikiText perplexity (lower is better).
Method Compression Zero-shot Accuracy (%) WikiText Perplexity
Phi-3 Mini Phi-3 Medium LLaMA-3 8B LLaMA-3 70B Phi-3 Mini Phi-3 Medium LLaMA-3 8B LLaMA-3 70B
Dense 0% 72.85 75.37 70.79 76.53 9.42 6.11 8.56 2.68
Uniform Sparsity Methods
SparseGPT 30% 70.63 74.53 69.08 75.07 11.19 7.48 9.71 3.24
40% 69.18 74.40 67.58 74.63 13.03 8.52 10.01 3.99
50% 66.36 73.25 64.66 73.17 16.80 9.89 11.95 5.27
Wanda 30% 70.66 74.05 68.63 75.19 10.71 7.28 9.39 3.28
40% 68.80 73.01 67.04 74.10 12.59 8.49 9.74 4.08
50% 65.03 70.96 63.27 72.85 17.23 10.12 12.36 5.38
DSNoT 30% 71.20 74.03 68.98 75.54 10.51 7.11 9.36 3.27
40% 69.08 72.90 66.65 74.29 12.17 8.24 9.60 4.10
50% 65.33 71.12 62.74 72.91 16.68 9.96 12.41 5.58
OATS 30% 71.48 74.04 69.34 75.24 10.27 6.85 9.59 3.07
40% 70.04 74.46 68.68 74.88 11.53 7.70 9.24 3.68
50% 68.41 73.39 65.71 73.30 15.18 9.05 10.87 4.78
Layerwise Allocation Methods (Based on Wanda)
OWL 30% 71.15 74.28 69.12 75.45 10.45 7.15 9.25 3.18
40% 69.32 73.35 67.58 74.42 12.28 8.32 9.58 3.95
50% 65.78 71.38 63.95 73.25 16.85 9.88 12.18 5.25
AlphaPruning 30% 71.28 74.35 69.25 75.52 10.38 7.08 9.18 3.15
40% 69.45 73.48 67.72 74.55 12.15 8.25 9.48 3.88
50% 65.95 71.52 64.12 73.42 16.72 9.78 12.05 5.18
Our Method
CAP 30% 72.15 74.85 70.25 76.02 9.88 6.58 9.05 2.95
40% 70.58 74.78 69.38 75.45 11.15 7.42 9.16 3.52
50% 69.12 74.05 66.85 74.18 14.68 8.78 10.35 4.45
@@ 3 Experiments
In this section, we first introduce the experimental setup. Subsequently, we present the main experimental results, extensions to modern instruction-tuned models, and detailed ablation studies. We further provide in-depth analyses on inference throughput and calibration robustness. Due to space constraints, detailed results for Llama-1/2 are provided in Appendix I, and thorough analyses of throughput performance and computational resource consumption of our proposed CAP method are presented in Appendix J.
@@
We evaluate our proposed CAP method on a comprehensive set of widely adopted large language models across different architectures and scales. Our evaluation includes the LLaMA family: LLaMA-1 (Touvron et al., 2023a) (7B, 13B, 30B), LLaMA-2 (Touvron et al., 2023b) (7B, 13B), and LLaMA-3 (Dubey et al., 2024) (8B, 70B); Modern Instruction-Tuned models: LLaMA-3.1-8B and Qwen2.5-7B (Yang et al., 2024); the OPT series (Zhang et al., 2022) (1.3B, 2.7B, 6.7B, 13B); the Phi-3 family (Abdin et al., 2024) including Phi-3 Mini (3.8B) and Phi-3 Medium (14B); and BERT-base (Devlin et al., 2019). To assess the performance of the compressed models, we conduct experiments on zero-shot tasks and language modeling. We perform an extensive evaluation of the zero-shot capabilities of pruned models across eight standard commonsense benchmark datasets: GLUE (Wang et al., 2019a), PIQA (Bisk et al., 2020), BoolQ (Clark et al., 2019), HellaSwag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2021), OpenBookQA (Mihaylov et al., 2018), and the ARC Easy and ARC Challenge tasks (Clark et al., 2018). For language modeling evaluation, we measure perplexity on the held-out WikiText (Merity et al., 2016) validation set (lm-eval-harness). For modern models, we additionally evaluate Chain-of-Thought reasoning (GSM8K (Bai et al., 2025)) and LongBench-v2 (Cobbe et al., 2021).
@@
We utilize PyTorch 2.3.0, Transformers 4.28.0, CUDA 12.1 on NVIDIA A100 GPUs under Ubuntu. To ensure fair comparison, we use 128 sequences with context length sampled from the C4 training set (Raffel et al., 2020) as calibration data. For policy gradient estimation, we set iterations to 3, sliding window size to 5, and learning rate to 0.05. The $\lambda$ parameter for RPCA decomposition is set according to the established formulation $\lambda=1/\sqrt{\max(m,n)}$, where $m$ and $n$ represent the dimensions of the data matrix. Further ablation studies on the $\lambda$ setting can be found in Appendix H.